# Trajectory Volatility for Out-of-Distribution Detection in Mathematical Reasoning

### v1

@@ A.4 Extended Conclusion
In Lemma A.2, we have proved the upper bound of differential-integral error order estimation is the equivalent infinitesimals of $1/L$. Actually, when $L\rightarrow+\infty$, we have
@@
Appendix B Motivation II: Empirical Analysis
B.1 Visualization Projection Setting
B.2 Visualization Example
Input Output
Domain 1: Algebra
Domain 2: Geometry
Domain 3: Number Theory
Domain 4: Precalculus
Simplify $\frac{1}{1-\tan^{2}x}+\frac{1}{1-\cot^{2}x}.$ 1
Input Output
Domain 1: Bible
Domain 2: News
Domain 3: TED
Domain 4: Health
Appendix C Algorithm: Pseudo-Code of TV Score Computation
@@
0:  $L$: The number of hidden layers $N$: The size of ID dataset $k$: the smoothing differential order of TV score $\{{\boldsymbol{y}_{l}}\}_{1\leq l\leq L}$: the average hidden embedding of the OOD sample in each layer $\{[\hat{{\boldsymbol{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$: the average hidden embedding of all $N$ ID samples in each layer
@@
2:     Fitting ID samples $\{[\hat{{\boldsymbol{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$ to Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l},{\boldsymbol{\Sigma}}_{l})$
3:     $f({\boldsymbol{y}_{l}})\leftarrow\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\boldsymbol{y}_{l}}-{\boldsymbol{\mu}_{l}}\right)$
@@
7:        $\Delta^{(i)}\mathcal{G}_{l}=\mathcal{N}({\boldsymbol{\mu}}_{l}^{(i)},{\boldsymbol{\Sigma}}_{l}^{(i)})\leftarrow\mathcal{N}(\sum_{t=0}^{i}(-1)^{i+t}\mathrm{C}_{i}^{t}{\boldsymbol{\mu}}_{l+i},~{}\sum_{t=0}^{i}\mathrm{C}_{i}^{t}{\boldsymbol{\Sigma}}_{l+i})$
@@
9:     $f^{(i)}({\boldsymbol{y}_{l}})\leftarrow\left({\boldsymbol{y}_{l}}^{(i)}-{\boldsymbol{\mu}_{l}}^{(i)}\right)^{\top}\left(\Sigma_{l}^{(i)}\right)^{-1}\left({\boldsymbol{y}_{l}}^{(i)}-{\boldsymbol{\mu}_{l}}^{(i)}\right)$
@@
12:     $\textrm{TV}_{i}\leftarrow\texttt{average}\left[f^{(i)}({\boldsymbol{y}_{l}})\right]_{1\leq l\leq L}$
@@
Appendix D Experimental Setting Details
D.1 Basic Information of Dataset
@@
Table 6: ID and OOD datasets used in this paper.
D.2 ID Dataset Split
@@ D.3 Training Implementation
D.4 OOD Dataset Selection Rationality
@@
D.5 Baseline
Let ${\boldsymbol{x}}$ represent the input sequence and ${\boldsymbol{y}}$ the output sequence, with lengths denoted as $n_{x}$ and $n_{y}$ respectively. In addition, we assume that $p(\cdot;{\boldsymbol{\theta}})$ represents the GLM parameterized by ${\boldsymbol{\theta}}$ that has been trained in ID dataset $\mathcal{D}$, outputting a sequence of softmax probabilities. We compare some training-free baseline methods as below:
@@ D.5 Baseline
    $\sum_{i=1}^{n_{y}}\max\frac{p(y_{i}|y_{\prec i},{\boldsymbol{x}};{\boldsymbol{\theta}})}{n_{y}}.$ (34)
@@
    $\int p({\boldsymbol{y}}|{\boldsymbol{x}};{\boldsymbol{\theta}})q({\boldsymbol{\theta}})\mathrm{d}\theta,$ (35)
    where $q({\boldsymbol{\theta}})\approx p({\boldsymbol{\theta}}|\mathcal{D})$.
@@
    $\left[\prod_{i=1}^{n_{y}}p(y_{i}|y_{\prec i},{\boldsymbol{x}};{\boldsymbol{\theta}})\right]^{-\frac{1}{N}}.$ (36)
@@
    $({\boldsymbol{x}}-{\boldsymbol{\mu}_{x}})^{\top}{\boldsymbol{\Sigma}_{x}^{-1}}({\boldsymbol{x}}-{\boldsymbol{\mu}_{x}}),$ (37)
    where ${\boldsymbol{\mu}_{x}}$ and ${\boldsymbol{\Sigma}_{x}}$ represent the mean and variance of the Gaussian distribution associated with the first-layer hidden state.
@@
    $({\boldsymbol{y}}-{\boldsymbol{\mu}_{y}})^{\top}{\boldsymbol{\Sigma}_{y}^{-1}}({\boldsymbol{y}}-{\boldsymbol{\mu}_{y}}),$ (38)
    where ${\boldsymbol{\mu}_{y}}$ and ${\boldsymbol{\Sigma}_{y}}$ represent the mean and variance of the Gaussian distribution associated with the final-layer hidden state.
Appendix E Full Main Experimental Results
OOD Detection.
AUROC and FPR95 results of each dataset are shown in Table 8 (Llama2-7B) and Table 9 (GPT2-XL).
OOD Quality Estimation.
Kendall’s $\tau$ correlation results of each dataset are shown in Table 10 (Llama2-7B) and Table 11 (GPT2-XL).
@@ Appendix E Full Main Experimental Results
MS-Prob [14] 79.97$\pm$1.60 / 80.97$\pm$4.90 82.60$\pm$1.25 / 82.57$\pm$4.62 63.87$\pm$1.64 / 96.19$\pm$1.57 76.20$\pm$1.46 / 85.20$\pm$2.69 90.68$\pm$0.94 / 62.27$\pm$4.03 78.66$\pm$1.38 / 81.44$\pm$3.56
MC-Drop [10] 72.12$\pm$1.68 / 85.43$\pm$6.60 74.23$\pm$1.74 / 81.17$\pm$6.28 55.64$\pm$2.87 / 97.15$\pm$1.34 67.53$\pm$2.08 / 89.09$\pm$5.41 73.61$\pm$2.66 / 82.36$\pm$4.78 68.63$\pm$2.21 / 87.04$\pm$4.88
PPL [3] 84.17$\pm$1.43 / 52.75$\pm$5.13 88.67$\pm$1.36 / 53.32$\pm$5.38 77.42$\pm$1.97 / 71.28$\pm$3.90 83.88$\pm$2.13 / 68.90$\pm$4.22 94.05$\pm$0.43 / 19.03$\pm$3.17 85.64$\pm$1.46 / 53.06$\pm$4.36
I-Emb [46] 81.51$\pm$1.09 / 62.84$\pm$4.56 75.41$\pm$0.97 / 69.66$\pm$2.07 62.53$\pm$1.30 / 88.35$\pm$1.72 84.42$\pm$0.85 / 54.80$\pm$7.25 75.59$\pm$0.92 / 63.71$\pm$2.87 75.89$\pm$1.03 / 67.87$\pm$3.69
O-Emb [46] 76.95$\pm$1.54 / 74.92$\pm$3.02 78.72$\pm$1.31 / 80.97$\pm$2.23 61.43$\pm$1.40 / 88.45$\pm$1.58 70.23$\pm$1.36 / 80.18$\pm$1.71 86.97$\pm$1.32 / 51.51$\pm$2.27 74.86$\pm$1.39 / 75.21$\pm$2.16
@@
MS-Prob [14] 53.08$\pm$1.67 / 94.07$\pm$1.86 56.56$\pm$1.53 / 90.06$\pm$2.39 63.31$\pm$1.88 / 87.36$\pm$2.42 66.68$\pm$1.32 / 86.15$\pm$2.61 61.07$\pm$1.30 / 86.91$\pm$2.76 60.14$\pm$1.54 / 88.91$\pm$2.41
MC-Drop [10] 48.87$\pm$2.43 / 96.78$\pm$1.21 44.90$\pm$2.76 / 92.33$\pm$2.09 57.34$\pm$1.57 / 89.15$\pm$2.34 54.09$\pm$2.42 / 90.07$\pm$1.96 56.46$\pm$1.85 / 91.21$\pm$1.83 52.33$\pm$2.21 / 91.92$\pm$1.89
PPL [3] 52.24$\pm$2.57 / 95.56$\pm$1.21 55.12$\pm$1.80 / 89.24$\pm$2.09 62.88$\pm$1.76 / 80.96$\pm$2.34 67.14$\pm$1.34 / 81.38$\pm$1.96 59.39$\pm$1.98 / 83.30$\pm$1.83 59.35$\pm$1.89 / 86.09$\pm$1.89
I-Emb [46] 45.68$\pm$1.50 / 95.05$\pm$1.17 60.92$\pm$1.34 / 86.97$\pm$4.13 66.28$\pm$0.92 / 76.09$\pm$2.93 61.18$\pm$1.22 / 87.36$\pm$1.89 67.60$\pm$1.20 / 77.78$\pm$2.53 60.33$\pm$1.37 / 84.65$\pm$2.53
O-Emb [46] 35.39$\pm$1.24 / 91.22$\pm$1.27 36.77$\pm$1.14 / 90.41$\pm$1.61 63.08$\pm$0.92 / 77.12$\pm$3.11 43.70$\pm$1.02 / 86.69$\pm$0.91 43.55$\pm$0.99 / 86.87$\pm$1.04 44.50$\pm$1.06 / 86.46$\pm$1.59
@@
Table 8: OOD Detection — Offline Discrimination (Llama2-7B): AUROC and FPR95 results ($p$-value > $0.05$ are grayed out). Underline denote SOTA among all baselines and bold denote SOTA among all methods.
@@
MS-Prob 72.13$\pm$1.42 / 78.35$\pm$2.35 64.09$\pm$1.78 / 87.42$\pm$1.24 68.45$\pm$1.56 / 82.35$\pm$1.57 69.37$\pm$1.43 / 82.09$\pm$1.90 78.67$\pm$0.89 / 61.23$\pm$3.02 70.54$\pm$1.42 / 78.29$\pm$2.02
MC-Drop 62.41$\pm$2.10 / 88.02$\pm$1.79 66.63$\pm$1.82 / 82.49$\pm$1.12 59.32$\pm$2.03 / 92.34$\pm$1.54 63.88$\pm$1.92 / 86.71$\pm$1.95 70.19$\pm$1.47 / 73.90$\pm$1.87 66.18$\pm$1.87 / 84.69$\pm$1.65
PPL 84.24$\pm$1.01 / 63.78$\pm$2.38 82.17$\pm$0.89 / 67.90$\pm$2.45 79.12$\pm$1.20 / 72.03$\pm$1.23 72.40$\pm$1.37 / 71.65$\pm$1.45 86.19$\pm$0.75 / 47.27$\pm$2.98 80.82$\pm$1.04 / 64.53$\pm$2.10
I-Emb 86.23$\pm$0.78 / 46.12$\pm$1.45 83.20$\pm$0.94 / 53.29$\pm$2.06 79.58$\pm$1.43 / 60.45$\pm$2.95 89.44$\pm$0.64 / 52.49$\pm$1.87 92.86$\pm$0.39 / 34.32$\pm$2.16 86.26$\pm$0.84 / 49.33$\pm$2.10
O-Emb 78.13$\pm$1.14 / 64.46$\pm$2.78 77.98$\pm$1.17 / 68.82$\pm$3.26 71.07$\pm$1.67 / 84.63$\pm$4.24 80.25$\pm$0.97 / 56.23$\pm$4.05 82.34$\pm$0.86 / 54.08$\pm$2.79 77.95$\pm$1.16 / 65.64$\pm$3.42
@@
MS-Prob 54.06$\pm$1.58 / 92.45$\pm$2.37 65.50$\pm$1.43 / 73.62$\pm$4.18 70.60$\pm$0.92 / 77.56$\pm$2.76 79.96$\pm$0.87 / 57.38$\pm$1.95 65.47$\pm$1.18 / 80.37$\pm$2.02 67.12$\pm$1.20 / 76.27$\pm$2.66
MC-Drop 63.44$\pm$2.08 / 79.85$\pm$3.35 62.13$\pm$1.65 / 77.28$\pm$3.21 67.97$\pm$1.62 / 71.42$\pm$2.88 71.29$\pm$1.43 / 69.22$\pm$2.41 52.89$\pm$1.80 / 92.78$\pm$0.67 63.54$\pm$1.72 / 78.08$\pm$2.50
PPL 72.89$\pm$1.56 / 74.63$\pm$1.52 70.79$\pm$1.38 / 79.65$\pm$1.21 67.50$\pm$1.04 / 87.76$\pm$0.92 87.14$\pm$0.57 / 43.09$\pm$1.23 70.40$\pm$1.06 / 76.80$\pm$1.45 73.74$\pm$1.12 / 72.39$\pm$1.27
I-Emb 87.58$\pm$1.14 / 48.45$\pm$2.06 84.81$\pm$1.20 / 62.75$\pm$5.68 80.34$\pm$0.64 / 46.62$\pm$2.17 81.44$\pm$0.49 / 49.05$\pm$3.24 81.91$\pm$0.93 / 57.65$\pm$2.67 83.22$\pm$0.88 / 52.90$\pm$3.16
O-Emb 82.05$\pm$1.36 / 60.44$\pm$3.96 81.42$\pm$2.02 / 68.43$\pm$4.38 72.98$\pm$0.85 / 66.72$\pm$1.77 80.00$\pm$1.02 / 70.23$\pm$1.87 79.94$\pm$0.94 / 57.68$\pm$1.60 79.28$\pm$1.24 / 64.70$\pm$2.72
@@
Table 9: OOD Detection — Offline Discrimination (GPT2-XL): AUROC and FPR95 results ($p$-value > $0.05$ are grayed out). Underline denote SOTA among all baselines and bold denote SOTA among all methods.

### v2

@@ C.4 Extended Conclusion
In Lemma C.2, we have proved the upper bound of differential-integral error order estimation is the equivalent infinitesimals of $1/L$. Actually, when $L\rightarrow+\infty$, we have
@@
Appendix D Algorithm: TV Score Computation
@@
0:  $L$: The number of hidden layers $N$: The size of ID dataset $k$: the smoothing differential order of TV score $\{{\bm{y}_{l}}\}_{1\leq l\leq L}$: the average hidden embedding of the OOD sample in each layer $\{[\hat{{\bm{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$: the average hidden embedding of all $N$ ID samples in each layer
@@
2:     Fitting ID samples $\{[\hat{{\bm{y}_{l}}}]_{i}\}_{1\leq l\leq L,1\leq i\leq N}$ to Gaussian distribution $\mathcal{G}_{l}=\mathcal{N}({\bm{\mu}}_{l},{\bm{\Sigma}}_{l})$
3:     $f({\bm{y}_{l}})\leftarrow\left({\bm{y}_{l}}-{\bm{\mu}_{l}}\right)^{\top}\left(\Sigma_{l}\right)^{-1}\left({\bm{y}_{l}}-{\bm{\mu}_{l}}\right)$
@@
7:        $\Delta^{(i)}\mathcal{G}_{l}=\mathcal{N}({\bm{\mu}}_{l}^{(i)},{\bm{\Sigma}}_{l}^{(i)})\leftarrow\mathcal{N}(\sum_{t=0}^{i}(-1)^{i+t}\mathrm{C}_{i}^{t}{\bm{\mu}}_{l+i},~{}\sum_{t=0}^{i}\mathrm{C}_{i}^{t}{\bm{\Sigma}}_{l+i})$
@@
9:     $f^{(i)}({\bm{y}_{l}})\leftarrow\left({\bm{y}_{l}}^{(i)}-{\bm{\mu}_{l}}^{(i)}\right)^{\top}\left(\Sigma_{l}^{(i)}\right)^{-1}\left({\bm{y}_{l}}^{(i)}-{\bm{\mu}_{l}}^{(i)}\right)$
@@
12:     $\textrm{TV}_{i}\leftarrow\texttt{average}\left[f^{(i)}({\bm{y}_{l}})\right]_{1\leq l\leq L}$
@@
As for the computational complexity, once we have all $\bm{y_{l}}$, our TV score requires only scalar addition and multiplication. During the ID distribution fitting phase, both operations are $\mathcal{O}(Ldn)$, where $n$ is the dataset size. During the score computation phase, both operations are $\mathcal{O}(Ldk)$. In practice, the computation time is measured in milliseconds.
Appendix E Experimental Setting Details
E.1 Basic Information of Dataset
@@
E.2 ID Dataset Split
@@ E.3 Training Implementation
E.4 OOD Dataset Selection Rationality
@@
E.5 Baseline
Let ${\bm{x}}$ represent the input sequence and ${\bm{y}}$ the output sequence, with lengths denoted as $n_{x}$ and $n_{y}$ respectively. In addition, we assume that $p(\cdot;{\bm{\theta}})$ represents the GLM parameterized by ${\bm{\theta}}$ that has been trained in ID dataset $\mathcal{D}$, outputting a sequence of softmax probabilities. We compare some training-free baseline methods as below:
@@ E.5 Baseline
    $\frac{1}{n_{y}}\cdot\sum_{i=1}^{n_{y}}p(y_{i}|y_{\prec i},{\bm{x}};{\bm{\theta}}).$ (30)
@@
    $\int p({\bm{y}}|{\bm{x}};{\bm{\theta}})q({\bm{\theta}})\mathrm{d}\theta,$ (31)
    where $q({\bm{\theta}})\approx p({\bm{\theta}}|\mathcal{D})$.
@@
    $\left[\prod_{i=1}^{n_{y}}p(y_{i}|y_{\prec i},{\bm{x}};{\bm{\theta}})\right]^{-\frac{1}{n_{y}}}.$ (32)
@@
    $({\bm{x}}-{\bm{\mu}_{x}})^{\top}{\bm{\Sigma}_{x}^{-1}}({\bm{x}}-{\bm{\mu}_{x}}),$ (33)
    where ${\bm{\mu}_{x}}$ and ${\bm{\Sigma}_{x}}$ represent the mean and variance of the Gaussian distribution associated with the first-layer hidden state.
@@
    $({\bm{y}}-{\bm{\mu}_{y}})^{\top}{\bm{\Sigma}_{y}^{-1}}({\bm{y}}-{\bm{\mu}_{y}}),$ (34)
    where ${\bm{\mu}_{y}}$ and ${\bm{\Sigma}_{y}}$ represent the mean and variance of the Gaussian distribution associated with the final-layer hidden state.
Appendix F Supplementary Experimental Results
F.1 Hyperparameter Ablation: Smoothing Order k
Figure 4: Smoothing order $k$ analysis: $k$ ranges from $0-5$ ($k=0$ corresponds to the original TV Score). The upper part is for the OOD detection scenario and the lower part is for the OOD quality estimation scenario; the left part is for the far-shift OOD datasets and the right part is for the near-shift OOD datasets.
F.2 OOD Detection
Results of each dataset are shown in Table 11 (Llama2-7B) and Table 12 (GPT2-XL).
Table 11: AUROC and FPR95 results in Llama2-7B of the Offline Detection scenario ($p$-value > $0.05$ are grayed out). Underline and bold denote SOTA among all baselines and all methods.
@@ F.2 OOD Detection
Max Softmax Prob. [12] 79.97$\pm$1.60 / 80.97$\pm$4.90 82.60$\pm$1.25 / 82.57$\pm$4.62 63.87$\pm$1.64 / 96.19$\pm$1.57 76.20$\pm$1.46 / 85.20$\pm$2.69 90.68$\pm$0.94 / 62.27$\pm$4.03 78.66$\pm$1.38 / 81.44$\pm$3.56
Monte-Carlo Dropout [8] 72.12$\pm$1.68 / 85.43$\pm$6.60 74.23$\pm$1.74 / 81.17$\pm$6.28 55.64$\pm$2.87 / 97.15$\pm$1.34 67.53$\pm$2.08 / 89.09$\pm$5.41 73.61$\pm$2.66 / 82.36$\pm$4.78 68.63$\pm$2.21 / 87.04$\pm$4.88
Perplexity [3] 84.17$\pm$1.43 / 52.75$\pm$5.13 88.67$\pm$1.36 / 53.32$\pm$5.38 77.42$\pm$1.97 / 71.28$\pm$3.90 83.88$\pm$2.13 / 68.90$\pm$4.22 94.05$\pm$0.43 / 19.03$\pm$3.17 85.64$\pm$1.46 / 53.06$\pm$4.36
Input Embedding [43] 81.51$\pm$1.09 / 62.84$\pm$4.56 75.41$\pm$0.97 / 69.66$\pm$2.07 62.53$\pm$1.30 / 88.35$\pm$1.72 84.42$\pm$0.85 / 54.80$\pm$7.25 75.59$\pm$0.92 / 63.71$\pm$2.87 75.89$\pm$1.03 / 67.87$\pm$3.69
Output Embedding [43] 76.95$\pm$1.54 / 74.92$\pm$3.02 78.72$\pm$1.31 / 80.97$\pm$2.23 61.43$\pm$1.40 / 88.45$\pm$1.58 70.23$\pm$1.36 / 80.18$\pm$1.71 86.97$\pm$1.32 / 51.51$\pm$2.27 74.86$\pm$1.39 / 75.21$\pm$2.16
@@
Max Softmax Prob. [12] 53.08$\pm$1.67 / 94.07$\pm$1.86 56.56$\pm$1.53 / 90.06$\pm$2.39 63.31$\pm$1.88 / 87.36$\pm$2.42 66.68$\pm$1.32 / 86.15$\pm$2.61 61.07$\pm$1.30 / 86.91$\pm$2.76 60.14$\pm$1.54 / 88.91$\pm$2.41
Monte-Carlo Dropout [8] 48.87$\pm$2.43 / 96.78$\pm$1.21 44.90$\pm$2.76 / 92.33$\pm$2.09 57.34$\pm$1.57 / 89.15$\pm$2.34 54.09$\pm$2.42 / 90.07$\pm$1.96 56.46$\pm$1.85 / 91.21$\pm$1.83 52.33$\pm$2.21 / 91.92$\pm$1.89
Perplexity [3] 52.24$\pm$2.57 / 95.56$\pm$1.21 55.12$\pm$1.80 / 89.24$\pm$2.09 62.88$\pm$1.76 / 80.96$\pm$2.34 67.14$\pm$1.34 / 81.38$\pm$1.96 59.39$\pm$1.98 / 83.30$\pm$1.83 59.35$\pm$1.89 / 86.09$\pm$1.89
Input Embedding [43] 45.68$\pm$1.50 / 95.05$\pm$1.17 60.92$\pm$1.34 / 86.97$\pm$4.13 66.28$\pm$0.92 / 76.09$\pm$2.93 61.18$\pm$1.22 / 87.36$\pm$1.89 67.60$\pm$1.20 / 77.78$\pm$2.53 60.33$\pm$1.37 / 84.65$\pm$2.53
Output Embedding [43] 35.39$\pm$1.24 / 91.22$\pm$1.27 36.77$\pm$1.14 / 90.41$\pm$1.61 63.08$\pm$0.92 / 77.12$\pm$3.11 43.70$\pm$1.02 / 86.69$\pm$0.91 43.55$\pm$0.99 / 86.87$\pm$1.04 44.50$\pm$1.06 / 86.46$\pm$1.59
@@
Table 12: AUROC and FPR95 results in GPT2-XL of the Offline Detection scenario ($p$-value > $0.05$ are grayed out). Underline and bold denote SOTA among all baselines and all methods.
@@
Max Softmax Prob. [12] 72.13$\pm$1.42 / 78.35$\pm$2.35 64.09$\pm$1.78 / 87.42$\pm$1.24 68.45$\pm$1.56 / 82.35$\pm$1.57 69.37$\pm$1.43 / 82.09$\pm$1.90 78.67$\pm$0.89 / 61.23$\pm$3.02 70.54$\pm$1.42 / 78.29$\pm$2.02
Monte-Carlo Dropout [8] 62.41$\pm$2.10 / 88.02$\pm$1.79 66.63$\pm$1.82 / 82.49$\pm$1.12 59.32$\pm$2.03 / 92.34$\pm$1.54 63.88$\pm$1.92 / 86.71$\pm$1.95 70.19$\pm$1.47 / 73.90$\pm$1.87 66.18$\pm$1.87 / 84.69$\pm$1.65
Perplexity [3] 84.24$\pm$1.01 / 63.78$\pm$2.38 82.17$\pm$0.89 / 67.90$\pm$2.45 79.12$\pm$1.20 / 72.03$\pm$1.23 72.40$\pm$1.37 / 71.65$\pm$1.45 86.19$\pm$0.75 / 47.27$\pm$2.98 80.82$\pm$1.04 / 64.53$\pm$2.10
Input Embedding [43] 86.23$\pm$0.78 / 46.12$\pm$1.45 83.20$\pm$0.94 / 53.29$\pm$2.06 79.58$\pm$1.43 / 60.45$\pm$2.95 89.44$\pm$0.64 / 52.49$\pm$1.87 92.86$\pm$0.39 / 34.32$\pm$2.16 86.26$\pm$0.84 / 49.33$\pm$2.10
Output Embedding [43] 78.13$\pm$1.14 / 64.46$\pm$2.78 77.98$\pm$1.17 / 68.82$\pm$3.26 71.07$\pm$1.67 / 84.63$\pm$4.24 80.25$\pm$0.97 / 56.23$\pm$4.05 82.34$\pm$0.86 / 54.08$\pm$2.79 77.95$\pm$1.16 / 65.64$\pm$3.42
@@
Max Softmax Prob. [12] 54.06$\pm$1.58 / 92.45$\pm$2.37 65.50$\pm$1.43 / 73.62$\pm$4.18 70.60$\pm$0.92 / 77.56$\pm$2.76 79.96$\pm$0.87 / 57.38$\pm$1.95 65.47$\pm$1.18 / 80.37$\pm$2.02 67.12$\pm$1.20 / 76.27$\pm$2.66
Monte-Carlo Dropout [8] 63.44$\pm$2.08 / 79.85$\pm$3.35 62.13$\pm$1.65 / 77.28$\pm$3.21 67.97$\pm$1.62 / 71.42$\pm$2.88 71.29$\pm$1.43 / 69.22$\pm$2.41 52.89$\pm$1.80 / 92.78$\pm$0.67 63.54$\pm$1.72 / 78.08$\pm$2.50
Perplexity [3] 72.89$\pm$1.56 / 74.63$\pm$1.52 70.79$\pm$1.38 / 79.65$\pm$1.21 67.50$\pm$1.04 / 87.76$\pm$0.92 87.14$\pm$0.57 / 43.09$\pm$1.23 70.40$\pm$1.06 / 76.80$\pm$1.45 73.74$\pm$1.12 / 72.39$\pm$1.27
Input Embedding [43] 87.58$\pm$1.14 / 48.45$\pm$2.06 84.81$\pm$1.20 / 62.75$\pm$5.68 80.34$\pm$0.64 / 46.62$\pm$2.17 81.44$\pm$0.49 / 49.05$\pm$3.24 81.91$\pm$0.93 / 57.65$\pm$2.67 83.22$\pm$0.88 / 52.90$\pm$3.16
Output Embedding [43] 82.05$\pm$1.36 / 60.44$\pm$3.96 81.42$\pm$2.02 / 68.43$\pm$4.38 72.98$\pm$0.85 / 66.72$\pm$1.77 80.00$\pm$1.02 / 70.23$\pm$1.87 79.94$\pm$0.94 / 57.68$\pm$1.60 79.28$\pm$1.24 / 64.70$\pm$2.72
@@
F.3 Beyond Detection: OOD Quality Estimation
Results of each dataset are shown in Table 13 (Llama2-7B) and Table 14 (GPT2-XL).